# TD-NeRF: Novel Truncated Depth Prior for Joint Camera Pose and Neural Radiance Field Optimization

### v1

@@ III Method
The overview of our proposed method TD-NeRF is illustrated in Fig. 2. We tackle a crucial challenge in the current NeRF [7] research, which involves simultaneously optimizing the neural radiance fields and camera pose without given camera poses in both indoor and outdoor scenes (Sec. III-A). We first introduce a lightweight pre-trained depth estimation network, DPT [29] to obtain depth priors. To fully exploit depth priors and gain more accurate depth geometry, we propose a ray sampling strategy based on the truncated normal distribution [30] and depth priors called Truncated Depth-Based Sampling (TDBS). Further, to speed up the convergence during training, prevent falling into local minima, and make the rendered depth more accurate, we propose the coarse-to-fine training strategy for TDBS (Sec. III-B). Moreover, to enhance robustness against depth noise, we introduce a Gaussian Point Constraint (GPC) that measures the distance between inter-frame point clouds (Sec. III-C.). Therefore, we effectively utilize the coarse depth map and seamlessly incorporate it into the synchronous training of NeRF and camera pose estimation.
@@ III-A Preliminary
It essentially models a mapping function that utilizes Multi-Layer Perceptron (MLP) layers, denoted as $F_{\theta}$, which maps 3D points $x$ and viewing directions $\mathbf{d}$ to color $c$ and volume density $\sigma$, i.e., $F_{\theta}(x,d)\xrightarrow{}(c,\sigma)$. The rendering process can be described as follows: 1) Given a sequence of images, the camera ray $r(t)=o+td\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ is cast into the scene, and a set of points xx is sampled along the ray; 2) The network $F_{\theta}$ is employed to estimate the density $\sigma$ and color $c$ of this set of points; 3) Volume rendering is used to integrate along the ray and obtain the color values $C(\mathbf{r})$ for synthesizing the final image. The entire integration process between $[t_{n},t_{f}]$ is modeled as:
@@
In the problem of pose-unknown NeRF, we simultaneously optimize the pose PP and parameters $\theta$ of NeRF. In existing methods, most approaches directly minimize the photometric error to obtain the final result. The mathematical expression is as follows:
@@ IV-B1 On Camera Pose Estimation
TABLE II: Quantitative evaluation of novel view synthesis and camera pose estimation on Tanks and Temples. We adopt the COLMAP pose as the ground truth. Unlike the dataset provided by NoPe-NeRF [16], to simulate the bandwidth-constrained scenario, the original data is sampled at 1 frame every 5 frames, and the frequency is reduced from 30 fps to 6 fps. ( †: indicates our re-implementation in 6 fps.)
@@
Our proposed method significantly outperforms other baselines in all metrics for camera pose estimation. NoPe-NeRF [16] is the state-of-the-art NeRF designed for pose-unknown novel view synthesis. NeRFmm [13] is one of the key benchmarks in the pose-unknown NeRF approach. We evaluated its performance on three datasets. For LLFF [34] and Tanks and Temples [35], where the ground truth pose is not available, we utilized the pose provided by COLMAP [18] as the reference. As shown in Tabs. I and II. Our method demonstrates a significant reduction in errors on the LLFF, achieving improvements of 44.8%, 66.4%, and 49.8% compared to the state-of-the-art methods. Similarly, on the Tanks and Temples, our method achieves reductions in errors of 8.88%, 5.27%, and 10.02%, respectively. To further validate the robustness of our approach, we conducted experiments on the BLEFF [13], which contains the ground truth of the camera pose. Our method outperforms the current state-of-the-art method by a considerable margin in terms of error reduction. We provide quantitative results and a qualitative visualization in Tab. III and Fig. 4, respectively. Our method achieves better quantitative and qualitative results than previous state-of-the-art methods. This experiment demonstrates that our method can optimize the camera pose simultaneously in both indoor and outdoor scenes with large motion changes.
@@ IV-B3 Effectiveness of coarse-to-fine TDBS
The third experiment is to evaluate the effectiveness of our sampling strategy and coarse-to-fine training strategy: the accuracy of depth estimation and camera pose estimation and faster training convergence. As shown in Fig. 6, our TDBS strategy generates more accurate depth geometry than the state-of-the-art method. Specifically, the full-stage TDBS performs better at capturing details in nearby objects, while the innovative coarse-to-fine TDBS reliably reconstructs depth maps with realistic details at any distance. Furthermore, coarse-to-fine TDBS can also achieve better results in terms of novel view synthesis. Moreover, the effectiveness of our coarse-to-fine training strategy is substantiated by quantitative experiments, as indicated in the results shown in Tab. V. Compared with the uniform and full-stage training strategies, the average errors of the coarse-to-fine training strategy in camera pose estimation are reduced by 73.3% and 46%, respectively. Additionally, convergence comparisons between TDBS and the original method, illustrated in Fig. 5, reveal that TDBS not only accelerates the convergence speed during training but also improves the accuracy of pose estimation. Most importantly, TDBS achieves the previously established minimal error within an estimated 1000 epochs, a fraction (one-tenth) of the epochs previously needed, while successfully avoiding local optima, thereby allowing for continued optimization. This advancement is a testament to the robustness of our strategy in elevating model training efficiency and precision.

### v2

@@ III Method
The overview of our proposed method TD-NeRF is illustrated in Fig. 2. We tackle a crucial challenge in the current NeRF [7] research, which involves simultaneously optimizing the neural radiance fields and camera pose without given camera poses in both indoor and outdoor scenes (Sec. III-A). Firstly, we utilize a lightweight pre-trained depth estimation network, DPT [29] to obtain depth priors. To fully exploit depth priors and gain more accurate depth geometry, we propose a ray sampling strategy based on the truncated normal distribution [30] and depth priors called Truncated Depth-Based Sampling (TDBS). Further, to speed up the convergence during training, prevent falling into local minima, and make the rendered depth more accurate, we propose the coarse-to-fine training strategy for TDBS (Sec. III-B). Moreover, to enhance robustness against depth noise, we introduce a Gaussian Point Constraint (GPC) that measures the distance between inter-frame point clouds (Sec. III-C.). Therefore, we effectively utilize the coarse depth map and seamlessly incorporate it into the synchronous training of NeRF and camera pose estimation.
@@ III-A Preliminary
It essentially models a mapping function that utilizes Multi-Layer Perceptron (MLP) layers, denoted as $F_{\theta}$, which maps 3D points $x$ and viewing directions $\mathbf{d}$ to color $c$ and volume density $\sigma$, i.e., $F_{\theta}(x,d)\xrightarrow{}(c,\sigma)$. The rendering process can be described as follows: 1) Given a sequence of images, the camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ is casting into the scene, and a set of points $x$ is sampled along the ray; 2) The network $F_{\theta}$ is employed to estimate the density $\sigma$ and color $c$ of this set of points; 3) Volume rendering is used to integrate along the ray and obtain the color values $C(\mathbf{r})$ for synthesizing the final image. The entire integration process between $[t_{n},t_{f}]$ is modeled as:
@@
In the problem of pose-unknown NeRF, we simultaneously optimize the pose $P$ and parameters $\theta$ of NeRF. In existing methods, most approaches directly minimize the photometric error to obtain the final result. The mathematical expression is as follows:
@@ IV-B1 On Camera Pose Estimation
TABLE II: Quantitative evaluation of novel view synthesis and camera pose estimation on Tanks and Temples (averaged over 8 scenes). We adopt the COLMAP pose as the ground truth. Unlike the dataset provided by NoPe-NeRF [16], to simulate the bandwidth-constrained scenario, the original data is sampled at 1 frame every 5 frames, and the frequency is reduced from 30 fps to 6 fps. ( †: indicates our re-implementation in 6 fps.)
@@
Our proposed method significantly outperforms other baselines in all metrics for camera pose estimation. NoPe-NeRF [16], a state-of-the-art NeRF designed for pose-unknown novel view synthesis, and NeRFmm [13], a key benchmark in pose-unknown NeRF approaches, were evaluated on three datasets. For LLFF [34] and Tanks and Temples [35], where the ground truth pose is not available, we utilized the pose provided by COLMAP [18] as the reference. As shown in Tabs. I and II. Our method demonstrates a significant reduction in errors on the LLFF, achieving improvements of 44.8%, 66.4%, and 49.8% compared to the state-of-the-art methods. Similarly, on the Tanks and Temples, our method achieves reductions in errors of 8.88%, 5.27%, and 10.02%, respectively. To further validate the robustness of our approach, we conducted experiments on the BLEFF [13], which contains the ground truth camera poses. Our method outperforms the current state-of-the-art method by a considerable margin in terms of error reduction. We provide quantitative results and a qualitative visualization in Tab. III and Fig. 4, respectively. Our method achieves better quantitative and qualitative results than previous state-of-the-art methods. This experiment demonstrates that our method can optimize the camera pose simultaneously in both indoor and outdoor scenes with large motion changes.
@@ IV-B3 Effectiveness of coarse-to-fine TDBS
The third experiment evaluates the effectiveness of our sampling strategy and coarse-to-fine training approach, focusing on depth estimation accuracy, camera pose estimation, and training convergence speed. As illustrated in Fig. 6, our TDBS strategy outperforms the state-of-the-art method in generating accurate depth geometry. Specifically, full-stage TDBS excels at capturing details in nearby objects, while the coarse-to-fine TDBS consistently reconstructs depth maps with realistic details across varying distances. Additionally, coarse-to-fine TDBS yields superior results in novel view synthesis. The effectiveness of our training strategy is further supported by quantitative results shown in Tab. V, where the coarse-to-fine approach reduces average camera pose estimation errors by 73.3% and 46% compared to uniform and full-stage strategies, respectively. Moreover, convergence analysis, depicted in Fig. 5, demonstrates that TDBS not only accelerates training convergence but also enhances pose estimation accuracy. Notably, TDBS achieves minimal error within approximately 1000 epochs—only one-tenth of the epochs previously required—while avoiding local optima and enabling continuous optimization. This highlights the robustness of our strategy in improving both training efficiency and model precision.